OpenStudy (anonymous):

Your Salary was 30200 in 2007 and 33500 in 2009. Your salary follows a linear growth pattern. What is salary in 2012? do i use y=3300x+30200 to get answer?

OpenStudy (gw2011):

You can use the following formula: Let x=The number of years needed y=30,200+1650x (1650 is the growth per year) y=30,200+1650(5) y=30,200+8250 y=38,450 This is the salaqry for 2012

OpenStudy (anonymous):

okay i see cause 3300 would be for two years thanks!!

OpenStudy (gw2011):

Right-you're welcome